OpenStudy (anonymous):

giving medal for answer What is the measure of each interior angle of a regular heptagon? Round to the nearest tenth. a. 96.2° c. 112.2° b. 98.6° d. 128.6°

OpenStudy (anonymous):

sum of interior angles formula for polygon is = 180(n-2) where n is the number of vértices. Heptagon n=7 so: 180(7-2)=180*5=900 to find the size of each angle just devide this by 7: 900/7=128.6°